OpenStudy (anonymous):

Solve for A: 2A/3 = 8 + 4A

OpenStudy (anonymous):

2A=3(8+4A) 2A=24+12A 2A-12A=24 -10A=24 A=24/-10 A= 2.4

OpenStudy (anonymous):

thanks for the help